* TT1 is in place, aligned so beam hits center of TT1, hits center of MMT1 (used pitch biases to finish pitch).
I had asked Q to write this down on a piece of paper, but then I forgot to transcribe it into the elog....
The TT screen matrix, at least for TT1, is flipped or something. When Eric moved the pit slider, the optic moved in yaw, and vice versa.
We need to fix this, but for now, here's the situation when TT1 was pointed correctly at MMT1:
TT1 Pit slider | 1000 1000 | ---> 700 UL
0 | -1000 1000 | ---> 700 LL
TT1 Yaw slider | 1000 -1000 | ---> -700 UR
0.7 | -1000 -1000 | ---> -700 LR
The confusing thing is that Koji and I confirmed (by plugging in the correct cable to the correct sensor) that "UL" on the screen goes to the UL coil, and the same for the other 3 coils. This needs investigation / fixing.